# Predictively Oriented Posteriors

### v1

@@ 3.2 Key Results
Equipped with the different forms of misspecification, we now compare the impact of different misspecification regimes on PrO posteriors $Q_{n}$, and compare them to Gibbs posteriors $Q_{n}^{\dagger}$. Whenever $Q_{n}$ and $Q_{n}^{\dagger}$ are compared directly, we assume that both are constructed from the same scoring rule. While Gibbs posteriors are always assumed to be constructed based on exact scoring rules, PrO posteriors may be based on approximations which adhere to $\operatorname{(DI)}$ or $\operatorname{(MS)}$ of Assumption 2.
We first present results applicable to exact and tractable predictive scores, and then obtain similar results for the setting of approximate predictive scores. Doing so reveals important differences in the approximation quality of approaches based on $\operatorname{(DI)}$ versus $\operatorname{(MS)}$, and allows us to obtain corollaries for PrO posteriors based on squared MMD as well as the logarithmic scoring rule. A full discussion of all underlying regularity conditions and a careful dissemination of the associated technical challenges is deferred to Section 7.
3.2.1 Tractable Predictive Scoring Rules
Before turning to the case of approximate predictive scores, it is informative to investigate the best case scenario: when the exact predictive scoring rule is tractable and (6) holds with equality, we obtain the strongest results.
@@ 3.2.1 Tractable Predictive Scoring Rules
Suppose that the predictive scoring rule is tractable so that (6) holds with equality. If Assumptions 13 and additional regularity conditions hold, then the following are satisfied for $n$ sufficiently large.
@@
    If the model is well-specified, $\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}},P_{0}\right)\right]-\mathbb{E}[\mathcal{D}_{S}(P_{Q_{n}^{\dagger}},P_{0})]\leq\dfrac{\log(n)}{n^{1/2}}$;
@@
    If the model is non-trivially misspecified, $\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}},P_{0}\right)\right]<\mathbb{E}[\mathcal{D}_{S}(P_{Q_{n}^{\dagger}},P_{0})]$.
    With further assumptions, $\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}},P_{0}\right)\right]\leq\inf_{Q\in\mathcal{P}(\Theta)}\mathcal{D}_{\mathcal{S}}(P_{Q},P_{0})+\dfrac{\log(n)}{n^{1/2}}$;
@@
    If the model allows convex recovery, then $\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}},P_{0}\right)\right]\leq\dfrac{\log(n)}{n^{1/2}}$.
[…]
The above result assumes that the exact predictive scoring rule is tractable, so that $\widehat{S}=S$ and (6) holds. Lemma 3 in Appendix B shows that this is satisfied for a number of scoring rules, for which Theorem 1 applies immediately.
Lemma 2.
For all kernel scores such as the squared MMD, the energy score, and the CRPS, the predictive scoring rule is tractable so that (6) holds exactly.
Interestingly, this result constitutes the only guarantees for the approach recently introduced in Shen et al. (2025), and shows that what was motivated as little more than a heuristic does in fact deliver strong and rigorous performance guarantees.
3.2.2 Approximate Scores
While the special case of exact predictive scores is very useful, certain scoring rules do not take the form in (6), and need to be approximated. To analyse what happens in this setting, one requires a refined version of Theorem 1 which introduces the difference between the exact and approximate score as $\operatorname{Gap}(k)$. While the below result denotes this difference generically, its magnitude and behaviour depend on the exact approximation chosen.
Theorem 2.
If Assumptions 13 and additional regularity conditions hold, then for the same $C_{k}$ as in Lemma 1, the following are satisfied for $n$ sufficiently large: for $n$ sufficiently large but finite, it holds for the same $C_{k}$ as in Lemma 1 that
  • $\operatorname{(WS)}$
    If the model is well-specified, $\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}},P_{0}\right)\right]-\mathbb{E}[\mathcal{D}_{S}(P_{Q_{n}^{\dagger}},P_{0})]\leq C_{k}\log(n)\sqrt{\frac{k}{n}}$;
  • $\operatorname{(NT)}$
    If the model is non-trivially misspecified, $\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}},P_{0}\right)\right]<\mathbb{E}[\mathcal{D}_{S}(P_{Q_{n}^{\dagger}},P_{0})]$.
    With further assumptions, $\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}},P_{0}\right)\right]\leq\inf_{Q\in\mathcal{P}(\Theta)}\mathcal{D}_{\mathcal{S}}(P_{Q},P_{0})+\operatorname{Gap}(k)+C_{k}\log(n)\sqrt{\frac{k}{n}}$;
  • $\operatorname{(CR)}$
    If the model allows convex recovery, then $\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}},P_{0}\right)\right]\leq\operatorname{Gap}(k)+C_{k}\log(n)\sqrt{\frac{k}{n}}$.
Intriguingly, the above shows that so long as predictive scoring rules are approximated in line with Assumption 2, our key conclusions regarding the predictive performance of PrO posteriors are the same as in Theorem 1: the PrO posterior performs comparably well to the Gibbs posterior in the well-specified setting, and strictly dominates it in the non-trivially misspecified setting. In fact, the $\operatorname{Gap}(k)$ term is only relevnt in situations where we bound the expected predictive performance of the PrO posterior directly, and without comparison to the Gibbs posterior. To tackle these bounds, we need to delineate along the axes of $\operatorname{(DI)}$ and $\operatorname{(MS)}$ defined in Assumption 2. The case of DI-based approximations to predictive scores is straightforward: since $k$ is fixed and cannot be varied, the gap term reverts to a fixed constant. In other words, relative to exact predictive scores, PrO posteriors based on inexact DI-based approximations may lead to an irreducible error of size $\operatorname{Gap}(k)$. In contrast, for MS-based PrO posteriors, $\operatorname{Gap}(k)\to 0$ as $k\to\infty$. To ensure that the bounds become non-vacuous however, the increase in $k$ needs to be carefully balanced against $n$ to prevent $C_{k}\log(n)\sqrt{\frac{k}{n}}$ from diverging. The rate at which $k$ can be allowed to increase will vary depending on the score chosen. For example, for the special case of the logarithmic scores, we obtain the following result.

### v2

@@ 3.2 Key Results
Having established the necessary regularity conditions, we proceed by studying the impact of different misspecification regimes on PrO posteriors $Q_{n}$, and compare them to Gibbs posteriors $Q_{n}^{\dagger}$ constructed using the same scoring rule. To state our results, we define
[…]
and take $\lambda_{n}\asymp{n}^{1/2}$ up to logarithmic factors. This allows us to state our first key result on the predictive performance of PrO posteriors.
@@
Assumption 1 is satisfied, and Assumption 2 or 2 holds. Under prior regularity conditions, the following holds for $n$ sufficiently large:
@@
    If the model is well-specified, $\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}},P_{0}\right)\right]-\mathbb{E}[\mathcal{D}_{S}(P_{Q_{n}^{\dagger}},P_{0})]\lesssim\nu_{n}$;
@@
    If the model is non-trivially misspecified, $\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}},P_{0}\right)\right]<\mathbb{E}[\mathcal{D}_{S}(P_{Q_{n}^{\dagger}},P_{0})]$,
    and $\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}},P_{0}\right)\right]-\inf_{Q\in\mathcal{P}(\Theta)}\mathcal{D}_{\mathcal{S}}(P_{Q},P_{0})\lesssim\nu_{n};$
@@
    If the model allows convex recovery, then $\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}},P_{0}\right)\right]\lesssim\nu_{n}.$
Theorem 1 evaluates the PrO posterior predictive $P_{Q_{n}}$ and its relationship to the Gibbs posterior predictive $P_{Q_{n}^{\dagger}}$ in terms of the divergence $\mathcal{D}_{S}$ induced by the score chosen to construct both posteriors. $\operatorname{(WS)}$ concerns the well-specified setting. Here, Gibbs posteriors typically concentrate towards the correct model parameter at the usual $\log(n)/n^{1/2}$ rate. As the result shows, the difference in predictive performance between PrO posteriors and Gibbs posteriors in this setting is small, and usually of order $\log(n)/n^{1/2}$. For the case of non-trivial misspecification studied in $\operatorname{(NT)}$, the situation changes: now, the PrO posteriors strictly outperforms Bayes and Gibbs posteriors. In particular, this performance gap will be of order $\inf_{\theta\in\Theta}\mathcal{D}_{S}(P_{\theta},P_{0})-\inf_{Q\in\mathcal{P}(\Theta)}\mathcal{D}_{S}(P_{Q},P_{0})$, which grows larger as the benefits of averaging over $\mathcal{M}_{\Theta}$ become more pronounced. To confirm this intuition, the result in $\operatorname{(CR)}$ considers the special case of convex recovery, where $\inf_{Q\in\mathcal{P}(\Theta)}\mathcal{D}_{S}(P_{Q},P_{0})=0$. As expected, the PrO posterior recovers the true data-generating process at rate $\nu_{n}$. In contrast, the Gibbs posterior predictive $P_{Q_{n}^{\dagger}}$ never recovers the true data generating process in this setting, and instead collapses onto a point mass at $\theta^{\star}=\operatornamewithlimits{argmin}_{\theta\in\Theta}\mathcal{D}_{S}(P_{\theta},P_{0})$. Lastly, the result also shows that the only case in which the choice of scoring rule has a notable impact is the non-trivially misspecified case: when the model satisfies convex recovery or is well-specified, any scoring rule will recover the same predictive distribution for $n$ sufficiently large. This suggests that PrO posteriors are generally less sensitive to the choice of scoring rule than Gibbs posteriors.
The above result validates the use of PrO posteriors whenever predictive performance matters, but sheds little light on the qualitative behaviour of PrO posteriors as expressions of parameter uncertainty. Given that they no longer collapse to a point mass, is it safe to assume that they behave well as the sample size increases? Indeed, in the context of parameter inference, convergence to a limiting distribution as more data are observed is fundamental to ensure that one is learning a well-defined target. Our second result assuages these concerns, and shows that PrO posteriors do indeed converge to a well-defined target given by the predictively optimal model average $Q^{*}:=\operatornamewithlimits{argmin}_{Q\in\mathcal{P}(\Theta)}\mathcal{D}_{S}(P_{Q},P_{0})$. Below, we state this convergence relative to a distance $\mathsf{d}$ defined on probability distributions and dependent on additional regularity conditions (see Section 8 for details).